Examine which of the following conicoids are central and which are non-central. Also determine which of the central conicoids have centre at the origin.

i) x2 +y2 +z2 + x + y + z = 1

ii) 2x2 +4xy+ xz−x−3y + 5z+ 3 = 0

iii) x2 −y2 −z2 +xy + 4yz + x = 0


1
Expert's answer
2021-06-01T16:29:38-0400

A conicoid given by equation ax2+by2+cz2+2fyz+2gxz+2hxy+2ux+2vy+2wz+d=0 has a point (x0,y0,z0) as a center if

"\\begin{cases}\n ax_ \n0\n\u200b\n +hy_0 \n\n\u200b\n +gz_ \n0\n\u200b\n +u=0 &\\ \\\\\n hx_ \n0\n\u200b\n +by_ \n0\n\u200b\n +fz_ \n0\n\u200b\n +v=0 \\\\\ngx_ \n0\n\u200b\n +fy_0 \n\n\u200b\n +cz_ \n0+w=0\\\\\n\\end{cases}"


1)

"a=b=c=1, u=v=w=1\/2, d=-1, f=g=h=0"

"\\begin{cases}\n x_ \n0\n\u200b\n +1\/2=0 \\\\\n y_ \n0\n\u200b\n +1\/2=0 \\\\\nz_ \n0\n\u200b\n +1\/2=0\\\\\n\\end{cases}"

This is the central conicoid with a center (-1/2,-1/2,-1/2)


2)

"a=2, h=2, g=1\/2, u=-1\/2, v=-3\/2, w=5\/2, d=3, b=c=f=0"

"\\begin{cases}\n 2x_0 \n\n\u200b\n +2y_0 \n\n\u200b\n +z_ \n0\n\u200b\n \/2\u22121\/2=0 \\\\\n 2x_ \n0\n\u200b\n \u22123\/2=0 \\\\\nx_ \n0\/2+\n\u200b\n \n\n\u200b\n 5\/2=0\\\\\n\\end{cases}"

This is not a central conicoid.


3)

"a=1, b=c=-1, h=1\/2, f=2, u=1\/2, d=g=v=w=0"

"\\begin{cases}\n x_ \n0\n\u200b\n +y_ \n0\n\u200b\n \/2+1\/2=0 \\\\\n x_ \n0\n\u200b\n \/2\u2212y_ \n0\n\u200b\n +2z_ \n0\n\u200b\n =0\\\\\n2y_0\n\u200b\n \n\n\u200b\n-z_0=0\\\\\n\\end{cases}"

"\\begin{cases}\n x_0 \n\n\u200b\n =\u22121\/2\u2212z_ \n0\n\u200b\n \/4 \\\\\n \u22121\/4\u2212z_ \n0\n\u200b\n \/8\u2212z_ \n0\n\u200b\n \/2+2z_ \n0\n\u200b\n =0\\\\\ny_0=z_0\/2\n\u200b\n \n\n\u200b\n\\\\\n\\end{cases}"

"\\begin{cases}\n x_0=-9\/11\n \\\\\n z_0=2\/11\n\u200b\n \n\u200b\n \\\\\ny_0=1\/11\n\u200b\n \n\n\u200b\n\\\\\n\\end{cases}"

This is the central conicoid with a center  (-6/11,1/11,2/11)